# AI-Driven Robotics for Free-Space Optics

### v1

@@ III Automated Optical Measurements
Figure 4: Automated optical measurements and characterization tasks. (a) The robotic arm linearly moves a camera, tracking the centroid of the laser beam to measure its direction and width. (b) The robotic arm places a lens in the beam path and moves the camera in the z-direction, capturing images. This data is used calculate the beam matrix. (c) The robotic arm assembles a setup consisting of a camera and a linear polarizer (LP) in the beam path. The beam passes through a birefringent film (BRF), creating spatial variations in the polarization. The robotic arm holds a quarter-wave plate (QWP) and rotates it, while the camera captures images. These images are used to calculate the Stokes parameters and plot polarization ellipses superimposed on the intensity image. (d) The robotic arm assembles a setup consisting of a grating and a camera to create a spectrometer and then holds and rotates a multilayer photonic crystal to obtain the angle- and wavelength-resolved transmission spectrum.
@@
Following the successful autonomous assembly and fine alignment of optical setups, we proceed to executing automated optical measurements with our platform. While the assembly process involves static placement of components, many optical experiments require dynamic operations such as translating, rotating, or adjusting optical elements in real time. These dynamic measurement tasks demand the same level of precision, robustness, and autonomy as the assembly process, but introduce additional challenges related to motion control. Moreover, measurements are also integral to the assembly phase itself—for instance, verifying laser alignment, and characterizing beam quality during setup construction. Thus, the integration of real-time optical diagnostic and measurement capabilities is critical to achieving fully autonomous experimentation.
@@
Leveraging the robotic platform’s ability to manipulate optical elements with high spatial and angular accuracy, we demonstrate automated execution of optical measurements. In this section, we show this functionality across four examples of distinct measurement tasks: (1) beam direction and waist tracking, (2) beam quality characterization through beam matrix measurement, (3) polarization mapping via spatially resolved Stokes parameter extraction, and (4) angle- and wavelength-resolved spectroscopy of multilayer photonic crystals.
@@
To determine the propagation direction of a collimated laser beam, the robotic arm picks up a camera module and positions it such that the beam is centered within the field of view. The robotic arm then translates the camera while keeping the beam centered in the camera’s field of view. By recording the beam centroid through this motion, the beam’s angle relative to the laboratory axes is calculated using a linear regression fit (see Supplementary Video 5). Similarly, the beam waist can be directly measured from the camera images. In a typical configuration, beam angles along the x, y, and z axes are measured to be ${}^{\circ}\pm$ 0.05, ${}^{\circ}\pm$ 0.05, and ${}^{\circ}\pm$ 0.04, respectively, averaged over 230 measurements (Figure 4a).
@@
Precise spatial characterization of laser beams is essential in a wide range of experiments, from beam shaping and mode matching to the calibration of complex optical systems. The complete spatial characterization of the beam can be determined via a measurement of the second-order moments of the Wigner distribution, commonly referred to as the beam matrix, which represent the beam’s spatial intensity and phase in both position and momentum (or angle) space. This matrix records beam properties such as width, divergence, astigmatism, and ellipticity. The components of the beam matrix can be measured by imaging the beam at different planes along the propagation axis and at different positions using using cylindrical and spherical lenses and a camera.
@@
Here, $x$ and $y$ are the spatial coordinates of the beam, and $\theta_{x}$ and $\theta_{y}$ are the angular divergences in the $x$ and $y$ directions, respectively. The diagonal elements describe the beam’s spread (e.g., $\langle x^{2}\rangle$ is the beam width in the $x$-direction), while the off-diagonal elements describe correlations between position and angle or between different spatial directions. The effective beam propagation ratio is measured to be $M_{\text{eff}}^{2}=\frac{4\pi}{\lambda}[\det M]^{1/4}=$ 1.07$\pm$0.03, where $\lambda=632.8$ nm is the wavelength of light.
We next turn our attention to the characterization of spatially varying polarization fields, which is essential for identifying polarization singularities and vectorial field distributions in beam optics applications. To show this functionality the robotic arm assembles a setup consisting of a linear polarizer and camera and then inserts and rotates a quarter-wave plate (QWP) along the beam path (see Supplementary Video 7). The intensity at each camera pixel $I(\theta)$ as a function of QWP angle $\theta$ follows:
@@
Finally, we demonstrate automated angle-resolved spectral characterization of multilayer photonic crystals, a task commonly performed to measure the bands of periodic photonic materials. This measurement typically require precise and repeatable angular sweeps, making them labor-intensive and sensitive to alignment errors when performed manually. The robotic arm first assembles a setup consisting of a diffraction grating and a camera positioned along the path of a supercontinuum (SC) laser beam. It then picks up a multilayer photonic crystal (consisting of alternating layers of Si and SiO2 with an embedded defect layer to form a cavity) housed in a 3D-printed casing. The arm then rotates the photonic crystal through a range of incidence angles (see Supplementary Video 8). At each angle, the transmission spectrum is recorded, yielding the complete wavelength- and angle-resolved transmission spectrum (Figure 4d) (see Methods for more details).
These four demonstrations collectively highlight the flexibility of robotic systems for executing high-precision optical measurements. Together, they establish that a robotic arm can serve as a versatile experimental tool in optics and photonics, effectively replacing a range of custom, purpose-built characterization instruments with a single reconfigurable platform.
@@ IV Discussion and conclusion
We have demonstrated the feasibility and effectiveness of an AI-driven robotic platform for optical setup design, assembly, alignment, and measurement. By integrating artificial intelligence, a robotic arm equipped with multimodal computer vision, and modular alignment tools, we have established the first complete automation pipeline that substantially reduces the need for manual intervention in optics laboratories.
For LLM-based optical setup generation, we have shown that our fine-tuning approach significantly outperforms conventional prompt-engineering-based or zero-shot methods with respect to physical validity, user compliance, and cost- and token-efficiency. This indicates that domain-specific fine-tuning, combined with structural constraints, can produce models capable of reasoning over physically grounded design spaces—a capability of increasing importance for AI-assisted experimentation.
The robotic assembly and alignment system addresses a persistent bottleneck in optics: the precise placement and alignment of free-space optical components. Our demonstration of automated optical measurements—including beam characterization, polarization mapping, and spectroscopy tasks—illustrates the breadth of tasks amenable to full automation. These examples represent generalizable workflows that can be extended to material characterization, imaging applications, and remote beam monitoring, particularly in hazardous environments such as high-power laser and X-ray facilities.
Looking ahead, several opportunities exist to further enhance this platform. One promising direction is the incorporation of reinforcement learning algorithms for setup alignment and real-time optimization under experimental conditions. This is particularly relevant in optics, where the reward signal can be sparse and highly intermittent during alignment, posing a challenge well-suited to reinforcement learning techniques. Another avenue is the integration of simulation-in-the-loop frameworks, where optical modeling tools inform both LLM-based setup generation and robotic execution. Finally, cloud-based deployment would enable distributed and democratized access to advanced optical experimentation.
@@ V Acknowledgments
We acknowledge support from the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions). S.C., R.K.S. and L.H. acknowledge funding from the MIT Undergraduate Research Opportunities Program (UROP). Z.C. acknowledges support from the Mathworks Fellowship. This work is also supported in part by the U. S. Army Research Office through the Institute for Soldier Nanotechnologies at MIT, under Collaborative Agreement Number W911NF-23-2-0121. The computations in this paper were partly run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University, We also acknowledge support of Parviz Tayebati.
@@ VI Author contributions
M.S., S.V. and S.Z.U. conceived the original idea with input from D.R.E.; S.Z.U., S.C., R.K.S., and S.V. developed the robotic platform; S.Z.U., S.C. and R.K.S. performed all experiments and measurements with input from S.V. and L.H.; Z.C. and S.V. generated the optical setup dataset, performed fine tuning of the LLMs and validation of generated setups. The manuscript was written by S.V. and S.Z.U. with inputs from all authors. M.S. supervised the project.
@@ VII Data Availability
The data and codes that support the plots within this paper and other findings of this work are available from the corresponding authors upon reasonable request. Correspondence and requests for materials should be addressed to S.Z.U. (suddin@mit.edu) and S.V. (svaidya1@mit.edu).
@@ X.1 Optical setup dataset generation
To construct a diverse and laboratory aware dataset of optical layouts, we develop a data augmentation pipeline centered around a library of canonical optical setups: Michelson interferometer, Mach-Zehnder interferometer, Hong-Ou-Mandel interferometer and a 4f optical system. Each setup is encoded as a list of component tuples—each specifying a component ID (e.g., mirror, beamsplitter), spatial coordinates (x, y in cm), and optical axis orientation (in degrees). We assume all components to have a fixed height in the z-direction. The first tuple with the component ID “000” is always a ray representing the incoming laser beam. The spatial coordinates within this tuple are those of any point lying on the ray and the orientation component is its angle with the x axis.
@@
Resulting setups were subjected to rigorous physical validation to ensure experimental feasibility. Validation criteria included: (1) all components must lie within the reachable circular workspace of a 61 cm radius; (2) no component may intersect the robot’s 15 cm radius circular footprint at the origin; (3) inter-component spacing must exceed individual component dimensions (7.5 cm × 6.4 cm) to avoid spatial overlap; and (4) internal beam paths must not intersect the robotic arm’s footprint, determined by constructing a complete graph from the component coordinates and checking for overlap. Setups failing any of these constraints were discarded.
@@ X.2 LLM Fine tuning
A consistent system prompt was employed throughout both the fine-tuning and inference stages to clearly define the task scope, specify the desired output format (structured lists of component tuples: ID, x, y, orientation), and impose fundamental constraints such as workspace boundaries and component non-overlap. Full details of the system prompt and dataset specifications are provided in the Supplementary Material.
@@
Our fine-tuning protocol involved two sequential stages: Structure Learning – In the first stage, the LLaMA3.1-8B-Instruct model, initialized with the system prompt, was fine-tuned on a dataset (described above) comprising pairs of auto-captioned natural language descriptors and their corresponding structured optical layouts. The primary goal of this stage was to enable the model to produce syntactically correct and physically plausible component arrangements, consistent with the basic spatial rules defined during dataset generation.
@@
User Request Understanding – Following the initial structure learning, the model underwent further fine-tuning using prompts representing more specific user requests, such as prescribed component positions or beam path characteristics (e.g., “Generate a Michelson interferometer setup with the incoming laser beam at 30 degrees relative to the x-axis”). While the target outputs remained structured component lists, the prompts introduced more nuanced constraints, improving the model’s ability to flexibly interpret and satisfy user-defined experimental requirements.
@@
To establish baseline performance benchmarks, we evaluated two contemporary large language models: GPT-4o and DeepSeek R1. For the GPT-4o model, we assessed performance under two distinct prompting conditions: (1) zero-shot prompting, providing the model solely with the user request and the required output format specification, and (2) few-shot prompting, augmenting the prompt with illustrative examples of both valid and invalid optical layouts. For the DeepSeek R1 model, we employed a few-shot Chain-of-Thought (CoT) prompting. This approach was structured to guide the model through step-by-step reasoning for optical layout construction, supported by representative examples. The accuracy of all evaluated approaches was measured against a newly generated test dataset, utilizing a generation temperature of $0.5$ to moderate output randomness. Further comparative analyses of model performance and detailed descriptions of the prompt engineering techniques are documented in the Supplementary Material.
X.3 Validation of optical setups

### v2

@@ III Automated Optical Measurements
[…]
@@
Following the successful autonomous assembly and fine alignment of optical setups, we proceed to execute automated optical measurements with our platform. While the assembly process involves static placement of components, many optical experiments require dynamic operations such as translating, rotating, or adjusting optical elements in real time. These dynamic measurement tasks demand the same level of precision, robustness, and autonomy as the assembly process, but introduce additional challenges related to motion control. Moreover, measurements are also integral to the assembly phase itself—for instance, verifying laser alignment and characterizing beam quality during setup construction. Thus, the integration of real-time optical diagnostics and measurement capabilities is critical to achieving fully autonomous experimentation.
@@
Leveraging the robotic platform’s ability to manipulate optical elements with high spatial and angular accuracy, we demonstrate automated execution of optical measurements. In this section, we show this functionality across five examples of distinct tasks: (1) beam direction and waist tracking, (2) beam quality characterization through beam matrix measurement, (3) polarization mapping via spatially resolved Stokes parameter extraction, (4) angle- and wavelength-resolved spectroscopy of multilayer photonic crystals, and (5) real-space optimization of a beam expansion setup.
@@
To determine the propagation direction of a collimated laser beam, the robotic arm picks up a camera module and positions it such that the beam is centered within the field of view. The robotic arm then translates the camera while keeping the beam centered in the camera’s field of view. By recording the beam centroid through this motion, the beam’s angle relative to the laboratory axes is calculated using a linear regression fit (see Supplementary Video 5). Similarly, the beam waist can be directly measured from the camera images (Figure 4a).
@@
Precise spatial characterization of laser beams is essential in a wide range of experiments, from beam shaping and mode matching to the calibration of complex optical systems. The complete spatial characterization of the beam can be determined via a measurement of the second-order moments of the Wigner distribution, commonly referred to as the beam matrix, which represent the beam’s spatial intensity and phase in both position and momentum (or angle) space. This matrix records beam properties such as width, divergence, astigmatism, and ellipticity. The components of the beam matrix can be measured using cylindrical and spherical lenses and a camera to image the beam at different planes along the propagation axis and at different positions.
@@
[…]
We next turn our attention to the characterization of spatially varying polarization fields, which is essential for identifying polarization singularities and vectorial field distributions in beam optics applications. To show this functionality, the robotic arm assembles a setup consisting of a linear polarizer and camera and then inserts and rotates a quarter-wave plate (QWP) along the beam path (see Supplementary Video 7). The intensity at each camera pixel $I(\theta)$ as a function of QWP angle $\theta$ follows:
@@
We next demonstrate automated angle-resolved spectral characterization of multilayer photonic crystals, a task commonly performed to measure the bands of periodic photonic materials. This measurement typically requires precise and repeatable angular sweeps, making it labor-intensive and sensitive to alignment errors when performed manually. The robotic arm first assembles a setup consisting of a diffraction grating and a camera positioned along the path of a supercontinuum (SC) laser beam. It then picks up a multilayer photonic crystal (consisting of alternating layers of Si and SiO2 with an embedded defect layer to form a cavity) housed in a 3D-printed casing. The arm then rotates the photonic crystal through a range of incidence angles (see Supplementary Video 8). At each angle, the transmission spectrum is recorded, yielding the complete wavelength- and angle-resolved transmission spectrum (Figure 4d) (see Methods for more details).
To demonstrate the robotic arm’s capability for real-space optimization, we construct a beam expander consisting of two convex lenses with focal lengths $f_{1}=60$ mm and $f_{2}=125$ mm. While the theoretical optimal separation is $f_{1}+f_{2}$, practical considerations such as lens positioning within housing assemblies and finite lens thickness necessitate experimental optimization to achieve the desired beam expansion ratio of $f_{2}/f_{1}$. To demonstrate this real-space optimization, the robotic arm first positions the first lens. An adaptive gradient descent algorithm then determines the optimal placement of the second lens by iteratively varying its coordinates and minimizing the error function defined as the absolute difference between the target and measured beam widths. Figure 4e shows the optimization error as a function of lens separation for two experimental runs, each initialized from positions displaced to either side of the theoretical optimum. Both optimization sequences converge to the correct lens position roughly within ten iterations, successfully achieving the target beam expansion ratio of $f_{2}/f_{1}=2.08$, as confirmed by the beam profile measurements shown for the original beam and the beams recorded before and after optimization for each run.
These five demonstrations collectively highlight the flexibility of robotic systems for executing high-precision optical measurements. Together, they establish that a robotic arm can serve as a versatile experimental tool in optics and photonics, effectively replacing a range of custom, purpose-built characterization instruments with a single reconfigurable platform.
@@ IV Discussion and conclusion
We have presented the first AI-driven robotic platform capable of designing, assembling, aligning, and operating free-space optical experiments, demonstrating that full-stack automation in optics is both accurate and technically achievable. By integrating artificial intelligence, a robotic arm guided by computer vision, and modular custom alignment tools, we have established an automation pipeline that substantially reduces manual intervention and unlocks a new regime of speed, reproducibility, and remote operability in optics.
For LLM-based optical setup generation, we have shown that our fine-tuning approach significantly outperforms conventional prompt-engineering-based or zero-shot methods with respect to physical validity, user compliance, and token-efficiency. This indicates that domain-specific fine-tuning, combined with structural constraints, can produce models capable of reasoning over physically grounded design spaces—a capability of increasing importance for AI-assisted experimentation across disciplines.
The robotic system addresses a persistent bottleneck in optics: the precise placement and alignment of free-space optical components. Our demonstration of automated optical measurements—including beam characterization, polarization mapping, spectroscopy, and optimization tasks—illustrates the breadth of tasks amenable to full automation. These examples represent generalizable workflows that can be extended to material characterization, imaging applications, and remote beam monitoring, particularly in hazardous environments such as high-power laser and X-ray facilities.
Looking ahead, several opportunities exist to further enhance this platform. One promising direction is the incorporation of reinforcement learning algorithms for setup alignment and continuous optimization under real-world non-ideal conditions [22, 23]. This is particularly relevant in optics, where the reward signal can be sparse and highly intermittent during alignment, posing a challenge that may be overcome with modern machine learning techniques. Another avenue is the integration of simulation-in-the-loop frameworks, where optical modeling tools inform both LLM-based setup generation and robotic execution, leading to self-optimizing workflows. Finally, remote operation and cloud-based optical laboratories will enable distributed and democratized access to advanced optical experimentation.
@@ V Acknowledgments
We acknowledge support from the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions). S.C., R.K.S. and L.H. acknowledge funding from the MIT Undergraduate Research Opportunities Program (UROP). Z.C. acknowledges support from the Mathworks Fellowship. This work is also supported in part by the U. S. Army Research Office through the Institute for Soldier Nanotechnologies at MIT, under Collaborative Agreement Number W911NF-23-2-0121. The computations in this paper were partly run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University. We acknowledge funding from MIT Generative AI Impact Consortium (MGAIC). We also acknowledge support from Parviz Tayebati. Copyediting of this manuscript was performed in part using ChatGPT models (versions 4o and 5). We also thank Dr. Paola Rebusco for critical reading and editing of the manuscript.
@@ VI Author contributions
M.S., S.V. and S.Z.U. conceived the original idea with input from D.R.E.; S.Z.U., S.V., S.C., and R.K.S. developed the robotic platform; S.Z.U., S.C. and R.K.S. performed all experiments and measurements with input from S.V. and L.H.; Z.C. and S.V. generated the optical setup dataset, performed fine tuning of the LLMs and validation of generated setups. The manuscript was written by S.V. and S.Z.U. with inputs from all authors. M.S. supervised the project.
@@ VII Data Availability
The codes that support the plots within this paper and other findings of this work are available online in [24, 25, 26, 27, 28]. Correspondence and requests for materials should be addressed to S.Z.U. (suddin@mit.edu) and S.V. (svaidya1@mit.edu).
@@ IX Supplementary Material: Videos
  • Video 10: Building a beam cleaning 4f system.
  • Video 11: Real-space optimization to create a $2\times$ beam expander.
@@ X.1 Optical setup dataset generation
To construct a diverse and laboratory-aware dataset of optical layouts, we developed a data augmentation pipeline centered around a library of canonical optical setups: Michelson interferometer, Mach-Zehnder interferometer, Hong-Ou-Mandel interferometer, and a 4f optical system. Each setup is encoded as a list of component tuples—each specifying a component ID (e.g., mirror, beam splitter), spatial coordinates (x, y in cm), and optical axis orientation (in degrees). We assumed all components to have a fixed height in the z-direction. The first tuple with the component ID “000” is always a ray representing the incoming laser beam. The spatial coordinates within this tuple are those of any point lying on the ray and the orientation component is its angle with the x axis.
@@
Resulting setups were subjected to rigorous physical validation to ensure experimental feasibility. Validation criteria included: (1) all components must lie within the reachable circular workspace of a 61 cm radius; (2) no component may intersect the robot’s 15 cm radius circular footprint at the origin; (3) inter-component spacing must exceed individual component dimensions (7.5 cm × 6.4 cm) to avoid spatial overlap; and (4) internal beam paths must not intersect the robotic arm’s footprint, determined by constructing a complete graph from the component coordinates and checking for overlap. Setups failing any of these constraints were discarded. The code for optical setup generation and validation is available in [24].
@@ X.2 LLM fine tuning
A consistent system prompt was employed throughout both the fine-tuning and inference stages to clearly define the task scope, specify the desired output format (structured lists of component tuples: ID, x, y, orientation), and impose fundamental constraints such as workspace boundaries and component non-overlap. Full details of the system prompt and dataset specifications are provided in [25].
X.2.1 Fine tuning protocol
Figure S1: Training curves for LLM accuracy. Accuracy (%), defined as the fraction of correct setups over total attempts, of the QuanTA fine-tuned LLaMA model as a function of training step index (scaled in units of $10^{5}$). Each point corresponds to an evaluation checkpoint. The three stages are: structure learning (Stage 0), user request understanding with three requests (Stage 1), user request understanding with random requests (Stage 2).
@@ X.2.1 Fine tuning protocol
Our fine-tuning protocol involved two sequential stages: Structure Learning – In the first stage, the LLaMA3.1-8B-Instruct model, initialized with the system prompt, was fine-tuned on the optical setup dataset (described above) comprising an equal but random mixture of pairs of auto-captioned natural language descriptors and their corresponding structured optical layouts for the four photonic setups. The primary goal of this stage was to enable the model to learn and internalize the knowledge behind the production of syntactically correct and physically plausible component arrangements, consistent with the basic spatial rules defined during dataset generation.
@@
User Request Understanding – Following the initial structure learning, the model underwent further fine-tuning using prompts representing more specific user requests, such as prescribed component positions or beam path characteristics (e.g., “Describe a compact Michelson interferometer optical system. The system should have a laser beam entering at 318.9 degrees.”). While the target outputs remained structured component lists, the prompts introduced more nuanced constraints, improving the model’s ability to flexibly interpret and satisfy user-defined experimental requirements. This stage is further split into two sessions. In the first session, we train on samples where the user always asks for a fixed number of three requests, allowing the model to understand user requests in a relatively simple format. We relax this condition in the second session, to further guide the model in interpreting flexible user requests.
To further improve the diversity of samples, parts of user prompts were generated independently using various templates, with varying precision requirements, language style, notation differences, etc. Together with the randomly generated setup coordinates and angles, this created an exponentially large data distribution, accurately mimicking possible real human inputs, and preventing the model from relying on specific patterns in the generated sample pairs (examples of input output pairs can be found in the [25]).
The training curves across all three stages are shown in Figure S1.
X.2.2 Hyperparameters
During fine-tuning, we chose QuanTA dimensions of 64-16-4, resulting in 143.13M training parameters (1.75% of the full model). We used the AdamW optimizer with a weight decay of 0.01 throughout the training. During the first stage, and the first session of the second stage, we used a peak learning rate of 0.00001, and during the second session of the second stage, we chose a peak learning rate of 0.000002. For each stage or session, we started the training with an initial linear warm-up, followed by a cosine decay. The first stage was trained for 349500 iterations. The first session of the second stage was trained for 288000 iterations, and the second session of the second stage was trained for 308000 iterations, with the checkpoint at 282000 iteration picked based on accuracy evaluation on a validation dataset independently generated from both the training set and the test set. We used a batch size of 8 throughout the training, and each stage or session was trained for only a single epoch, with no samples seen twice by the model. The full training required $\sim$8500 GPU hours with A100 GPUs and 80GB VRAM on each GPU.
X.2.3 Other prompting strategies
@@
To establish baseline performance benchmarks, we evaluated two contemporary large language models: GPT-4o (May 13, 2024 version) and DeepSeek-R1 (Jan 20, 2025 version). For the GPT-4o model, we assessed performance under two distinct prompting conditions: (1) zero-shot prompting, providing the model solely with the user request and the required output format specification, with minimal examples for demonstrating the output format (the same prompt as the one used during fine-tuning of the LLaMA3.1-8B-Instruct model) and (2) few-shot prompting, augmenting the prompt with illustrative examples of valid pairs of user input and model output. For the DeepSeek-R1 model, we employed the same few-shot prompt, while allowing the model to perform a few-shot Chain-of-Thought (CoT) reasoning. This approach was structured to guide the model through step-by-step reasoning for optical layout construction, supported by representative examples.
X.3 Validation of optical setups and LLM performance